OpenStudy (anonymous):

Given f(x)=x^2+1,fg(x)=x^2+4x+5 Find the function of g. @mathmath333

OpenStudy (anonymous):

\[x^2(g(x))+1=x^2+4x+5\]

OpenStudy (anonymous):

separate your 5 \[fog(x)=(x^2+4x+4)+1\]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

\[fog(x)=f(g(x))\]It's a way of representing composite functions All I did was separate your 5 into 4 and 1 \[fog(x)=f(g(x)=x^2+4x+4+1\]

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

Now we must substitute f(x)=x^2+1 ?

OpenStudy (anonymous):

that is not needed really just notice your first 3 terms, they form an identity

OpenStudy (michele_laino):

pleasenote that we can write: \[f\left( {g\left( x \right)} \right) = {\left( {g\left( x \right)} \right)^2} + 1\]

OpenStudy (michele_laino):

furthermore, we have: \[{x^2} + 4x + 5 = {x^2} + 4x + 4 + 1 = {\left( {x + 2} \right)^2} + 1\]

OpenStudy (michele_laino):

nemely: \[f\left( {g\left( x \right)} \right) = {x^2} + 4x + 5 = {x^2} + 4x + 4 + 1 = {\left( {x + 2} \right)^2} + 1\]

OpenStudy (anonymous):

so the function of g is g(x)=x+2?

OpenStudy (michele_laino):

that's right!

OpenStudy (anonymous):

Thnx @Michele_Laino and Nishant_Garg

OpenStudy (michele_laino):

thanks! @MARC_